# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ Abstract
When training large models, such as neural networks, the full derivatives of order 2 and beyond are usually inaccessible, due to their computational cost. This is why, among the second-order optimization methods, it is very common to bypass the computation of the Hessian by using first-order information, such as the gradient of the parameters (e.g., quasi-Newton methods) or the activations (e.g., K-FAC).
In this paper, we focus on the exact and explicit computation of projections of the Hessian and higher-order derivatives on well-chosen subspaces, which are relevant for optimization. Namely, for a given partition of the set of parameters, it is possible to compute tensors which can be seen as “higher-order derivatives according to the partition”, at a reasonable cost as long as the number of subsets of the partition remains small.
Then, we propose an optimization method exploiting these tensors at order 2 and 3 with several interesting properties, including: it outputs a learning rate per subset of parameters, which can be used for hyperparameter tuning; it takes into account long-range interactions between the layers of the trained neural network, which is usually not the case in similar methods (e.g., K-FAC); the trajectory of the optimization is invariant under affine layer-wise reparameterization.
Code available at https://github.com/p-wol/GroupedNewton/.
@@ 1 Introduction
The appealing theoretical properties of Newton’s method have led to numerous attempts to adapt it to neural network optimization. Therefore, the study of the Hessian of a loss according to many parameters has become a research area in itself, leading to a large number of methods to approximate it accurately with low computational cost.
Newton’s method applied to neural network optimization.
When it comes to neural networks, Newton’s method suffers from several problems. Some are technical, such as building an accurate and computationally efficient method for estimating the Hessian. But some of them are essential, in the sense that they cannot be solved by a perfect knowledge of the full Hessian alone. For instance, several works (Sagun et al., 2018) have shown that many eigenvalues of the Hessian are close to zero when training neural networks, making it impossible to use Newton’s method in practice, even if the Hessian is perfectly known.
Therefore, in this work, we do not aim to build a technique to estimate the Hessian accurately and efficiently. Instead, we propose to take a step aside and focus on two related goals. First, we aim to efficiently access higher-order information in order to use it for optimization. To do this, we leverage the available computational tools, that is, automatic differentiation and partition of the set of parameters into tensors stored on a GPU. Second, we aim to build a second-order optimization method with properties similar to those of Newton’s method (but inevitably weaker). In particular, we want a method that is invariant by layer-wise affine reparameterizations of the model, and that provides a Hessian-inspired matrix showing the interactions between each pair of tensors, along with layer-wise step sizes. Naturally, the computational cost should remain reasonable, and the problem of close-to-zero eigenvalues of the Hessian has to solved.
First contribution: extracting higher-order information.
Formally, we study a loss $\mathcal{L}$ to minimize according to a vector of parameters $\bm{\theta}\in\mathbb{R}^{P}$, which can be represented as a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. In a multilayer perceptron with $L$ layers, the $S=2L$ tensors $(\mathbf{T}^{s})_{1\leq s\leq 2L}$ are the tensors of weights and the vectors of biases of each layer. In that case, $S\ll P$. Within this framework, we propose a technique summarizing the order-$d$ derivative of the loss, which is a tensor belonging to $\mathbb{R}^{P^{d}}$, into a tensor belonging to $\mathbb{R}^{S^{d}}$, which is significantly smaller and easier to compute.
Second contribution: second-order optimization method.
@@
Then, we make use of the preceding technique at orders $2$ and $3$ to build a second-order optimization method. Formally, the method presented here and Newton’s method look alike: in both cases, a linear system $\mathbf{H}_{0}\mathbf{x}=\mathbf{g}_{0}$ has to be solved (according to $\mathbf{x}$), where $\mathbf{g}_{0}$ and $\mathbf{H}_{0}$ contain respectively first-order and second-order information about $\mathcal{L}$. Despite this formal resemblance, the difference is enormous: with Newton’s method, $\mathbf{H}_{0}$ is equal to the Hessian $\mathbf{H}$ of $\mathcal{L}$ of size $P\times P$, while with ours, $\mathbf{H}_{0}$ is equal to a matrix $\bar{\mathbf{H}}$ of size $S\times S$. Thus, $\bar{\mathbf{H}}$ is undoubtedly smaller and easier to compute than $\mathbf{H}$ when $S\ll P$. Nevertheless, since $\bar{\mathbf{H}}$ is a dense matrix, it still contains information about the interactions between the tensors $\mathbf{T}^{s}$ when they are used in $\mathcal{L}$. This point is crucial because most second-order optimization methods applied to neural networks use a simplified version of the Hessian (or its inverse), usually a diagonal or block-diagonal approximation, which ignores the interactions between layers. Finally, we propose an anisotropic version of Nesterov’s cubic regularization (Nesterov & Polyak, 2006), which uses order-3 information to regularize $\bar{\mathbf{H}}$ and avoid instabilities when computing $\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}$.
A proof-of-concept of this method is given in Section 5. Apart from that, this method provides a way to evaluate the interactions between layers from a training perspective, it outputs by-layer learning rates (which makes it applicable in hyperparameter tuning), and its training trajectory is invariant by layer-wise affine reparameterizations (so it preserves some interesting theoretical properties of Newton’s method).
@@
First, we show the context and motivation of our work in Section 2. Then, we provide in Sections 3 and 4 standalone presentations of the contributions, respectively the higher-order information extraction technique and the second-order optimization method. In Section 5, we present experimental results showing that the developed methods are usable in practice. Finally, we discuss the results in Section 6.
@@ 2.1 Higher-order information
It is not a novel idea to extract higher-order information about a loss at a minimal computational cost to improve optimization. This is typically what is done by Dangel (2023), although it does not go beyond the second-order derivative. In this line of research, the Hessian-vector product (Pearlmutter, 1994) is a decisive tool, that allows to compute the projection of higher-order derivatives in given directions at low cost (see App. A). For derivatives of order $3$ and beyond, Nesterov’s cubic regularization of Newton’s method (Nesterov & Polyak, 2006) uses information of order $3$ to avoid too large training steps. Incidentally, we develop an anisotropic variant of this in Section 4.2.
2.2 Second-order methods
The Hessian $\mathbf{H}$ of the loss $\mathcal{L}$ according to the vector of parameters $\bm{\theta}$ is known to contain useful information about $\mathcal{L}$. Above all, the Hessian is used to develop second-order optimization algorithms. Let us denote by $\bm{\theta}_{t}$ the value of $\bm{\theta}$ at time step $t$, $\mathbf{g}_{t}\in\mathbb{R}^{P}$ the gradient of $\mathcal{L}$ at step $t$ and $\mathbf{H}_{t}$ its Hessian at step $t$. One of the most widely known second-order optimization method is Newton’s method, whose step is (Nocedal & Wright, 1999, Chap. 3.3):
$\bm{\theta}_{t+1}:=\bm{\theta}_{t}-\mathbf{H}_{t}^{-1}\mathbf{g}_{t}.$ (1)
Under certain conditions, including strong convexity of $\mathcal{L}$, the convergence rate of Newton’s method is quadratic (Nocedal & Wright, 1999, Th. 3.7), which makes it very appealing. Besides, other methods use second-order information without requiring the full computation of the Hessian. For instance, Cauchy’s steepest descent (Cauchy, 1847) is a variation of the usual gradient descent, where the step size is tuned by extracting very little information from the Hessian:
$\bm{\theta}_{t+1}:=\bm{\theta}_{t}-\eta_{t}^{*}\mathbf{g}_{t},\quad\text{where}\quad\eta_{t}^{*}:=\frac{\mathbf{g}_{t}^{T}\mathbf{g}_{t}}{\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}},$ (2)
where the value of $\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}$ can be obtained with little computational cost (see Appendix A). However, when optimizing a quadratic function $f$ with Cauchy’s steepest descent, $f(\theta_{t})$ is known to decrease at a rate $(\frac{\lambda_{\mathrm{max}}-\lambda_{\mathrm{min}}}{\lambda_{\mathrm{max}}+\lambda_{\mathrm{min}}})^{2}$, where $\lambda_{\mathrm{max}}$ and $\lambda_{\mathrm{min}}$ are respectively the largest and the smallest eigenvalues of the Hessian of $f$ (Luenberger & Ye, 2008, Chap. 8.2, Th. 2). If the Hessian of $f$ is strongly anisotropic, then this rate is close to one and optimization is slow. For a comparison of the two methods, see (Gill et al., 1981; Luenberger & Ye, 2008; Nocedal & Wright, 1999).
@@ 2.2 Second-order methods
When the parameter space is high-dimensional, computation of the Hessian $\mathbf{H}_{t}$ and inversion of the linear system $\mathbf{g}_{t}=\mathbf{H}_{t}\mathbf{x}$ are computationally intensive. Quasi-Newton methods are designed to avoid any direct computation of the Hessian, and make extensive use of gradients and finite difference methods to approximate the direction of $\mathbf{H}_{t}^{-1}\mathbf{g}_{t}$. For a list of quasi-Newton methods, see (Nocedal & Wright, 1999, Chap. 8). However, Nocedal & Wright (1999) argue that, since it is easy to compute the Hessian by using Automatic Differentiation (AutoDiff), quasi-Newton methods tend to lose their interest. Nevertheless, they should remain useful when such computation is too difficult.
@@
Many methods overcome the curse of the number of parameters by exploiting the structure of the neural networks. It is then common to neglect interactions between layers, leading to a (block)-diagonal approximation of the Hessian. A first attempt has been made by Wang & Lin (1998): they divide the Hessian into blocks, following the division of the network into layers, and its off-diagonal blocks are removed. From another perspective, Ollivier (2015) keeps this block-diagonal structure, but performs an additional approximation on the remaining blocks.
More recently, K-BFGS has been proposed (Goldfarb et al., 2020), which is a variation of the quasi-Newton method BFGS with block-diagonal approximation and an approximate representation of these blocks. In a similar spirit, the Natural Gradient method TNT (Ren & Goldfarb, 2021) also exploits the structure of neural networks by performing a block-diagonal approximation. Finally, AdaHessian (Yao et al., 2021) efficiently implements a second-order method by approximating the Hessian by its diagonal.
Kronecker-Factored Approximate Curvature (K-FAC).
This approximation of the Hessian was proposed in (Martens & Grosse, 2015) in the context of neural network training. K-FAC exploits the specific architecture of neural networks to output a cheap approximation of the true Hessian. Despite its scalability, K-FAC suffers from several problems. First, the main approximation is quite rough, since “[it assumes] statistical independence between products […] of unit activities and products […] of unit input derivatives” (Martens & Grosse, 2015, Sec. 3.1). Second, even with an approximation of the Hessian, one has to invert it, which is computationally intensive even for small networks. To overcome this difficulty, a block-(tri)diagonal approximation of the inverse of the Hessian is made, which eliminates many of the interactions between the layers.
@@
Also, (Lu et al., 2018) proposes to approximate the Hessian with a matrix composed of blocks in which all coefficients are identical. Thus, the Hessian can be compressed into a smaller matrix that looks like the summary of the Hessian matrix $\bar{\mathbf{H}}$ used in Section 4. Sketching the Hessian in an optimization has been proposed in Gower et al. (2019), which is driven by the same spirit as the method presented in Section 4: instead of dealing with a complicated large matrix, one should “project” it onto spaces of lower dimension. Such a method can also be completed with Nesterov cubic regularization Hanzely et al. (2020). In a completely different setup, Yuan et al. (2022) proposed a “Sketched Newton-Raphson”, which is a sketched Newton method for root-finding. Finally, the technique proposed in Nonomura et al. (2021) is also close to the one we propose, and the formalism used is almost the same, although the use-case is completely different.
@@
Several optimization methods, such as Newton’s, have an optimization step invariant by affine reparameterization of $\bm{\theta}$ (Amari, 1998) (Nesterov, 2003, Chap. 4.1.2). Specifically, when using Newton’s method, it is equivalent to optimize $\mathcal{L}$ according to $\bm{\theta}$ and according to $\tilde{\bm{\theta}}=\mathbf{A}\bm{\theta}+\mathbf{B}$ ($\mathbf{A}\in\mathbb{R}^{P\times P}$ invertible, $\mathbf{B}\in\mathbb{R}^{P}$). This affine-invariance property holds even if the function $\mathcal{L}$ to minimize is a negative $\log$-likelihood, and one chooses to minimize $\bm{\theta}$ by the natural gradient method (Amari, 1998). This method also requires computing the Hessian of $\mathcal{L}$ at some point.
@@
Finally, many methods acquire information about the curvature of the loss surface by using only the gradients. For instance, Shampoo (Gupta et al., 2018) uses second-moment information of the accumulated gradients.
@@ 2.3 Motivation
The methods that aim to estimate the Hessian matrix $\mathbf{H}$ or its inverse $\mathbf{H}^{-1}$ in order to imitate Newton’s method implicitly assume that Newton’s method is adapted to the current problem. This assumption is certainly correct when the loss to optimize is strongly convex. But, when the loss is not convex and very complicated, e.g. when training a neural network, this assumption is not justified. Worse, it has been shown empirically that, at the end of the training of a neural network, the eigenvalues of the Hessian are concentrated around zero (Sagun et al., 2018), with only a few large positive eigenvalues. Therefore, Newton’s method itself does not seem to be recommended for neural network training, so we may not need to compute the full Hessian at all, which would relieve us of a tedious, if not impossible, task.
To avoid such problems, it is very common to regularize the Hessian by adding a small, constant term $\lambda\mathbf{I}$ to it (Nocedal & Wright, 1999, Chap. 6.3). Also, trust-region Newton methods are designed to handle non-positive-definite Hessian matrices (Nocedal & Wright, 1999, Chap. 6.4) (Nash, 1984).
@@
Also, some empirical works have shown that the role and the behavior of each layer must be considered along its interactions with the other layers, which emphasize the importance of off-diagonal blocks in the Hessian or its inverse. We give two examples. First, Zhang et al. (2022) has shown that, at the end of their training, many networks exhibit a strange feature: some (but not all) layers can be reinitialized to their initial value with little loss of the performance. Second, Kornblith et al. (2019) has compared the similarity between the representations of the data after each layer: changing the number of layers can qualitatively change the similarity matrix of the layers (Kornblith et al., 2019, Fig. 3). Among all, these results motivate our search for mathematical objects that show how layers interact.
@@
A whole line of research is concerned with building a well-founded method for finding a good scaling for the initialization distribution of the parameters, and for the learning rates, which can be chosen layer-wise. For instance, a layer-wise scaling for the weights was proposed and theoretically justified in the paper introducing the Neural Tangent Kernels (Jacot et al., 2018). Also, in the “feature learning” line of work, (Yang & Hu, 2021) proposes a relationship between different scalings related to weight initialization and training. Therefore, there is an interest in finding a scalable and theoretically grounded method to build per-layer learning rates.

### v4

# Gathering and Exploiting Higher-Order Information when Training Large Structured Models

@@ Abstract
When training large models, such as neural networks, the full derivatives of order 2 and beyond are usually inaccessible, due to their computational cost. Therefore, among the second-order optimization methods, it is common to bypass the computation of the Hessian by using first-order information, such as the gradient of the parameters (e.g., quasi-Newton methods) or the activations (e.g., K-FAC).
In this paper, we focus on the exact and explicit computation of projections of the Hessian and higher-order derivatives on well-chosen subspaces relevant for optimization. Namely, for a given partition of the set of parameters, we compute tensors that can be seen as “higher-order derivatives according to the partition”, at a reasonable cost as long as the number of subsets of the partition remains small.
Then, we give some examples of how these tensors can be used. First, we show how to compute a learning rate per subset of parameters, which can be used for hyperparameter tuning. Second, we show how to use these tensors at order 2 to construct an optimization method that uses information contained in the Hessian. Third, we show how to use these tensors at order 3 (information contained in the third derivative of the loss) to regularize this optimization method. The resulting training step has several interesting properties, including: it takes into account long-range interactions between the layers of the trained neural network, which is usually not the case in similar methods (e.g., K-FAC); the trajectory of the optimization is invariant under affine layer-wise reparameterization.
@@ 1 Introduction
In machine learning, computing the derivatives of the loss at various orders is challenging when using large models, such as neural networks. While the first-order derivative is relatively cheap to compute and easy to use to train neural networks, things get difficult when it comes to higher-order derivatives. In particular, Hessian-based training algorithms such as Newton’s method are very expensive to use on large models. Therefore, the study of the Hessian of a loss according to many parameters has become a research area in its own right. For derivatives of order 3 and higher, the situation is even worse: their exact computation is far more expensive than the Hessian, and only a few optimization algorithms use them.
Main contribution: extracting higher-order information.
Formally, we study a loss $\mathcal{L}$ to be minimized according to a vector of parameters $\boldsymbol{\theta}\in\mathbb{R}^{P}$. Thus, the order-$d$ derivative of $\mathcal{L}$ at a given point is a tensor of order $d$ with $P^{d}$ coefficients. Usually, such tensors cannot be computed exactly and explicitly with $d\geq 2$ (which includes the Hessian) for medium-sized models ($P\gtrsim 10^{6}$).
Instead of trying to approximate these tensors, we propose to compute their projections along well-chosen directions, that are relevant for optimization. This computation can be done efficiently by taking advantage of the practical implementation of the vector of parameters $\boldsymbol{\theta}$ as a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. Such a projection of the order-$d$ derivative yields a tensor of order $d$ with $S^{d}$ coefficients, instead of $P^{d}$. Thus, the Hessian of a model of size $P=10^{6}$ represented by a tuple of $S=20$ tensors can be reduced to a matrix of size $S^{2}=400$, instead of $P^{2}=10^{12}$. More generally, whenever $S\ll P$, the projected order-$d$ derivative of $\mathcal{L}$ is significantly smaller and easier to compute than the full order-$d$ derivative.
Application: computing per-layer learning rates.
Then, we show that such projections of the order-1 and order-2 derivatives can be used to compute the optimal learning rates to choose for each one of the $S$ tensors (or subsets of parameters). The procedure we propose to compute per-layer learning rates is both theoretically well-grounded and usable in practice (as long as the number of layers is not too large). In particular, our computation does not neglect long-range interactions between layers.
Application: second-order optimization method.
@@
Finally, we show that the information contained in the 1st, 2nd, and 3rd order derivatives is not only accessible at reasonable cost, but can also be used for optimization. In particular, we propose an optimization method that exploits higher-order information about the loss obtained by using the main contribution. For simplicity, our optimization method and Newton’s method look similar: in both cases, a linear system $\mathbf{H}_{0}\mathbf{x}=\mathbf{g}_{0}$ has to be solved (w.r.t. $\mathbf{x}$), where $\mathbf{g}_{0}$ and $\mathbf{H}_{0}$ contain respectively first-order and second-order information about $\mathcal{L}$. Despite this formal resemblance, the difference is enormous: with Newton’s method, $\mathbf{H}_{0}$ is equal to the Hessian $\mathbf{H}$ of $\mathcal{L}$ of size $P\times P$, while with ours, $\mathbf{H}_{0}$ is equal to a matrix $\bar{\mathbf{H}}$ of size $S\times S$. Thus, $\bar{\mathbf{H}}$ is undoubtedly smaller and easier to compute than $\mathbf{H}$ when $S\ll P$. Nevertheless, since $\bar{\mathbf{H}}$ is a dense matrix, it still contains information about the interactions between the tensors $\mathbf{T}^{s}$ when they are used in $\mathcal{L}$. This point is crucial because most second-order optimization methods applied to neural networks use a simplified version of the Hessian (or its inverse), usually a diagonal or block-diagonal approximation, ignoring interactions between layers. Additionally, we propose an anisotropic version of Nesterov’s cubic regularization [24], which uses order-3 information to regularize $\bar{\mathbf{H}}$ and avoid instabilities when computing $\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}$. In particular, the resulting training trajectory is invariant by layer-wise affine reparameterizations, so our method preserves some interesting properties of Newton’s method.
@@
First, we show the context and motivation of our work in Section 2. Then, we provide in Section 3 our core method, and in Sections 4 and 5 its applications. In Section 6, we present experimental results showing that the developed methods are usable in practice. Finally, we discuss the results in Section 7.
@@ 2.1 Higher-order information
It is not a novel idea to extract higher-order information about a loss at a minimal computational cost to improve optimization. This is typically what is done by [6], although it does not go beyond the second-order derivative. In this line of research, the Hessian-vector product [28] is a decisive tool, that allows to compute the projection of higher-order derivatives in given directions at low cost (see App. 0.A). For derivatives of order $3$, Nesterov’s cubic regularization of Newton’s method [24] uses information of order $3$ to avoid too large training steps. Incidentally, we develop an anisotropic variant of this in Section 5. In the same spirit, the use of derivatives of any order for optimization has been proposed [3]
2.2 Using and estimating the Hessian in optimization
The Hessian $\mathbf{H}$ of the loss $\mathcal{L}$ according to the vector of parameters $\boldsymbol{\theta}$ is known to contain useful information about $\mathcal{L}$. Above all, the Hessian is used to develop second-order optimization algorithms. Let us denote by $\boldsymbol{\theta}_{t}$ the value of $\boldsymbol{\theta}$ at time step $t$, $\mathbf{g}_{t}\in\mathbb{R}^{P}$ the gradient of $\mathcal{L}$ at step $t$ and $\mathbf{H}_{t}$ its Hessian at step $t$. One of the most widely known second-order optimization method is Newton’s method, whose step is [25, Chap. 3.3]:
$\boldsymbol{\theta}_{t+1}:=\boldsymbol{\theta}_{t}-\mathbf{H}_{t}^{-1}\mathbf{g}_{t}.$ (1)
Under certain conditions, including strong convexity of $\mathcal{L}$, the convergence rate of Newton’s method is quadratic [25, Th. 3.7], which makes it very appealing. Besides, other methods use second-order information without requiring the full computation of the Hessian. For instance, Cauchy’s steepest descent [4] is a variation of the usual gradient descent, where the step size is tuned by extracting very little information from the Hessian:
$\boldsymbol{\theta}_{t+1}:=\boldsymbol{\theta}_{t}-\eta_{t}^{*}\mathbf{g}_{t},\quad\text{where}\quad\eta_{t}^{*}:=\frac{\mathbf{g}_{t}^{T}\mathbf{g}_{t}}{\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}},$ (2)
where the value of $\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}$ can be obtained with little computational cost (see Appendix 0.A). However, when optimizing a quadratic function $f$ with Cauchy’s steepest descent, $f(\theta_{t})$ is known to decrease at a rate $(\frac{\lambda_{\mathrm{max}}-\lambda_{\mathrm{min}}}{\lambda_{\mathrm{max}}+\lambda_{\mathrm{min}}})^{2}$, where $\lambda_{\mathrm{max}}$ and $\lambda_{\mathrm{min}}$ are respectively the largest and the smallest eigenvalues of the Hessian of $f$ [19, Chap. 8.2, Th. 2]. If the Hessian of $f$ is strongly anisotropic, then this rate is close to one and optimization is slow. For a comparison of the two methods, see [9, 19, 25].
@@ 2.2 Using and estimating the Hessian in optimization
When the parameter space is high-dimensional, computation of the Hessian $\mathbf{H}_{t}$ and inversion of the linear system $\mathbf{g}_{t}=\mathbf{H}_{t}\mathbf{x}$ are computationally intensive. Quasi-Newton methods are designed to avoid any direct computation of the Hessian, and make extensive use of gradients and finite difference methods to approximate the direction of $\mathbf{H}_{t}^{-1}\mathbf{g}_{t}$. For a list of quasi-Newton methods, see [25, Chap. 8]. However, [25] argue that, since it is easy to compute the Hessian by using Automatic Differentiation (AutoDiff), quasi-Newton methods tend to lose their interest.
@@
Many methods overcome the curse of the number of parameters by exploiting the structure of the neural networks. It is then common to neglect interactions between layers, leading to a (block)-diagonal approximation of the Hessian. A first attempt has been made by [32]: they divide the Hessian into blocks, following the division of the network into layers, and its off-diagonal blocks are removed. From another perspective, [27] keeps this block-diagonal structure, but performs an additional approximation on the remaining blocks.
More recently, K-BFGS has been proposed [10], which is a variation of the quasi-Newton method BFGS with block-diagonal approximation and an approximate representation of these blocks. In a similar spirit, the Natural Gradient method TNT [29] also exploits the structure of neural networks by performing a block-diagonal approximation. Finally, AdaHessian [34] efficiently implements a second-order method by approximating the Hessian by its diagonal.
Kronecker-Factored Approximate Curvature (K-FAC) is a method for approximating of the Hessian proposed in [20] in the context of neural network training. K-FAC exploits the specific architecture of neural networks to output a cheap approximation of the true Hessian. Despite its scalability, K-FAC suffers from several problems. First, the main approximation is quite rough, since “[it assumes] statistical independence between products […] of unit activities and products […] of unit input derivatives” [20, Sec. 3.1]. Second, even with an approximation of the Hessian, one has to invert it, which is computationally intensive even for small networks. To overcome this difficulty, a block-(tri)diagonal approximation of the inverse of the Hessian is made, which eliminates many of the interactions between the layers.
@@
In Section 5, we propose to summarize the Hessian to avoid the expensive computation of the full Hessian. This idea is not new. For instance, [18] proposes to approximate the Hessian with a matrix composed of blocks in which all coefficients are identical. A more broadly used technique to compress the Hessian is to perform sketching on it, that is, project it on randomly chosen directions. This idea is used for solving solve linear systems [35], as well as for minimizing functions [11], and can be further adapted to Newton’s method with cubic regularization [13]. Finally, it is also possible to choose the directions of the projection by using available information [26]. This is the strategy that we have adopted in Section 5.
@@
Several optimization methods, such as Newton’s, have an optimization step invariant by affine reparameterization of $\boldsymbol{\theta}$ [1] [23, Chap. 4.1.2]. Specifically, when using Newton’s method, it is equivalent to optimize $\mathcal{L}$ according to $\boldsymbol{\theta}$ and according to $\tilde{\boldsymbol{\theta}}=\mathbf{A}\boldsymbol{\theta}+\mathbf{B}$ ($\mathbf{A}\in\mathbb{R}^{P\times P}$ invertible, $\mathbf{B}\in\mathbb{R}^{P}$). This affine-invariance property holds even if the function $\mathcal{L}$ to minimize is a negative $\log$-likelihood, and one chooses to minimize $\boldsymbol{\theta}$ by the natural gradient method [1]. This method also requires computing the Hessian of $\mathcal{L}$ at some point.
@@
Finally, many methods acquire geometric information on the loss by using only the gradients. For instance, Shampoo [12] uses second-moment information of the accumulated gradients.
@@ 2.3 Motivation
The methods that aim to estimate the Hessian matrix $\mathbf{H}$ or its inverse $\mathbf{H}^{-1}$ in order to imitate Newton’s method implicitly assume that Newton’s method is adapted to the current problem. This assumption is certainly correct when the loss to optimize is strongly convex. But, when the loss is not convex and very complicated, e.g. when training a neural network, this assumption is not justified. Worse, it has been shown empirically that, at the end of the training of a neural network, the eigenvalues of the Hessian are concentrated around zero [30], with only a few large positive eigenvalues. Therefore, Newton’s method itself does not seem to be recommended for neural network training, so we may not need to compute the full Hessian at all, which would relieve us of a tedious, if not impossible, task.
To avoid such problems, it is very common to regularize the Hessian by adding a small, constant term $\lambda\mathbf{I}$ to it [25, Chap. 6.3]. Also, trust-region Newton methods are designed to handle non-positive-definite Hessian matrices [25, Chap. 6.4] [22].
@@
Also, some empirical works have shown that the role and the behavior of each layer must be considered along its interactions with the other layers, which emphasize the importance of off-diagonal blocks in the Hessian or its inverse. We give two examples. First, [36] has shown that, at the end of their training, many networks exhibit a strange feature: some (but not all) layers can be reinitialized to their initial value with little loss of the performance. Second, [15] has compared the similarity between the representations of the data after each layer: changing the number of layers can qualitatively change the similarity matrix of the layers [15, Fig. 3]. Among all, these results motivate our search for mathematical objects that show how layers interact.
@@
A whole line of research is concerned with building a well-founded method for finding a good scaling for the initialization distribution of the parameters, and for the learning rates, which can be chosen layer-wise. For instance, a layer-wise scaling for the weights was proposed and theoretically justified in the paper introducing the Neural Tangent Kernels [14]. Also, in the “feature learning” line of work, [33] proposes a relationship between different scalings related to weight initialization and training. Therefore, there is an interest in finding a scalable and theoretically grounded method to build per-layer learning rates.